Status Report

NOAA SATOPS Morning Report: Thursday, May 24, 2012

By SpaceRef Editor
May 24, 2012
Filed under ,

SATOPS Morning Report: Thursday, May 24, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

DOY 145

GOES
GOES-12 Imager instrument Cycle slips caused intermittent image degradation at DOY144 10:40z, DOY145 08:31z.

GOES-12 Wallops frame breaks due to BER 13:51z – 14:33z, 00:09z.
GOES-13 Wallops frame breaks due to BER 12:34z; 17:16z – 17:27z.
GOES-15 Wallops frame breaks due to BER 10:19z – 10:27z; 19:42z – 20:23z; 22:34z.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
Issues with POGO-B resolved.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
N -15 HRPT REV. 72942 AND REV. 72946 NOT RECEIVED AT ESPC.

ESPC IS NOT RECEIVING THE HRPT (HIGH RESOLUTION PICTURE TRANSMISSION) DATA FROM THE MONTEREY DIRECT READ-OUT SITE. MONTEREY IS INVESTIGATING.
DATE/TIME OF INITIAL IMPACT: MAY 22, 2012, 2100 UTC.

NCEP’s GFS model major upgrade and associated output product change to the GFS model output caused outages in ASCAT products (May 22, 2012 2100z to May 23, 2012 1402z; ~17 hours) and High Density Winds products (May 22, 2012 2100z to May 23, 2012 2200z; ~25 hours). All products are now recovered.

Questions or comments to [email protected]

SpaceRef staff editor.